# Causal Inference with High-dimensional Discrete Covariates

### v1

@@ 4 Role of Effect Homogeneity
In this section, we study the role of effect homogeneity in consistent estimation of treatment effects in the regime $n\lesssim d$. Let $\tau_{k}=\mathbb{E}[Y|X=k,A=1]-\mathbb{E}[Y|X=k,A=0]=\mu_{1k}-\mu_{0k}$ be the conditional average treatment effect (CATE) in the $k$-th category. In the high-dimensional regime where $d=d_{n}\rightarrow\infty$, $\tau_{k}$’s should be viewed as a triangular array $\{\tau_{nk},1\leq k\leq d_{n},n\geq 1\}$ and we will slightly abuse the notation to denote $\tau_{k}=\tau_{nk}$. Homogeneous effects (i.e. when $\tau_{k}=\psi$ for all $k\in[d]$) can be helpful in terms of estimation. Intuitively, since the treatment effects are the same within each level of covariate, there is no need to consistently estimate CATE $\tau_{k}$ for all possible $d$ categories. One only needs to estimate the CATE within a few categories accurately and due to effect homogeneity, these estimated CATEs generalize to other levels of covariate, which potentially reduces the sample size required for consistent estimation. In this work, we adopt a novel form of approximate effect homogeneity via parameter capturing the extent of heterogeneity. Mathematically, denote
@@
where $\widehat{t}_{k}=I(\widehat{p}_{k}>0,0<\widehat{\pi}_{k}<1)$ is the indicator of whether both treated and untreated samples are observed in the $k$-th level and $\widehat{\tau}_{k}=\widehat{\mu}_{1k}-\widehat{\mu}_{0k}$. The idea is to only estimate the CATE $\tau_{k}$ within those categories with both treated and untreated units and take the average over such categories. We restrict our attention to these categories so that unbiased estimation of $(\mu_{1k},\mu_{0k})$ (and hence $\tau_{k}$) is possible. We again define $/0=0$ 0 if $\sum_{k=1}^{d}\widehat{t}_{k}=0$, i.e. there is no such category that contains both treated and untreated units. Clearly on the event $\left\{\sum_{k=1}^{d}\widehat{t}_{k}=0\right\}$, we cannot obtain any information on treatment effects from $\widehat{\tau}$. Under appropriate scaling conditions, the probability of this “adverse” event converges to 00 even in the case $n\lesssim d$, as summarized in the following lemma.
@@
$\operatorname{Var}(\widehat{\tau})\leq\mathbb{E}\left[\operatorname{Var}\left(\widehat{\tau}\big{|}\mathbf{X}^{n},\mathbf{A}^{n}\right)\right]+\sigma_{n}^{2}+2\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right)+4\sigma_{n}\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right).$
If we further assume that there exists a constant $\alpha\in(0,1]$ such that $p_{k}\geq\alpha/d$ for all $k\in[d]$, then in the regime $n\leq\frac{3\epsilon(1-\epsilon)d}{\alpha}$, we have
$\mathbb{E}\left[\operatorname{Var}\left(\widehat{\tau}\big{|}\mathbf{X}^{n},\mathbf{A}^{n}\right)\right]\leq\frac{4d}{\alpha^{2}\epsilon(1-\epsilon)n^{2}}+\frac{3}{2}\exp\left(-\frac{\alpha^{2}\epsilon(1-\epsilon)n^{2}}{2d}\right)+\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right).$
Hence the estimator is consistent if $\sigma_{n}\rightarrow 0$ and $d/n^{2}\rightarrow 0$ in the regime $n\lesssim d$.
The assumption $p_{k}\geq\alpha/d$ is required in our proof on bounding the expected conditional variance. We allow $p_{k}$ to approach zero but the rate should be slower than $1/d$. An important special case is $\alpha=1$ which corresponds to the uniform distribution on $[d]$. It is common to impose some constraints on the proportion probability of each level in categorical data analysis (Zeng et al., 2023a, ; Gu and Xu,, 2023). In the literature of regression adjustment with high-dimensional covariates, this assumption may also be required. For instance, Lei and Ding, (2021) considered regression adjustment estimation of sample average treatment effects and they impose conditions on the maximum leverage score of the projection matrix $\mathbf{H}=\mathbf{X}(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top}$, where $\mathbf{X}$ is the design matrix. Their condition translates to lower bounds on the proportions $p_{k}$’s when the covariates are discrete. Furthermore, another plausible condition that requires $p_{k}$’s are comparable
$\delta\leq p_{i}/p_{j}\leq\delta^{-1},\,i\neq j,$
with $\delta\in(0,1)$ being a constant implies $p_{k}\gtrsim 1/d$. Such conditions can be frequently found in the high-dimensional testing literature (Jiang and Qi,, 2015; He et al.,, 2021).
Theorem 3 shows that if asymptotic effect homogeneity holds (i.e. $\sigma_{n}\rightarrow 0$), then the estimator $\widehat{\tau}$ is consistent in the regime $d/n^{2}\rightarrow 0$. Compared with $d/n\rightarrow 0$ required for consistency by the plug-in estimator $\widehat{\psi}$, $\widehat{\tau}$ has a faster convergence rate and enables us to achieve consistency in a wider regime under asymptotic effect homogeneity.
@@ 5.1 Average Treatment Effects
The negative result in Theorem 4 can be explained by the fact that the underlying measure of ATE estimation is the product of covariate distribution and propensity score, as shown in analyzing its second-order estimator (Robins et al.,, 2009; Zeng et al., 2023b, ). In observational studies where propensity score is unknown, it’s hard to obtain nearly unbiased estimates of $p_{k}\mu_{1k}$ (the summand in $\psi_{1}$) with the observed data. One may write $p_{k}\mu_{1k}=p_{k}\pi_{k}\mu_{1k}/\pi_{k}$ with $p_{k}\pi_{k}\mu_{1k}$ unbiasedly estimable. The extra term $1/\pi_{k}$ could be approximated by a polynomial with possibly diverging degrees
@@ B.6 Proof of Lemma 1
Lemma 5.
(Theorem 5.10 in Mitzenmacher and Upfal, (2017)) Let $\mathbf{X}^{n}\in\mathbb{R}^{d}\sim\text{Multinomial}\left(n,p_{1},\ldots,p_{d}\right)$, $\mathbf{Y}^{n}\in\mathbb{R}^{d}$ and $Y_{i}^{n}\sim\text{Poisson}(np_{i})$ and $Y_{1}^{n},\dots,Y_{d}^{n}$ are independent. Consider a non-negative function $f(x_{1},\dots,x_{d})$, if $\mathbb{E}[f(X_{1}^{n},\dots,X_{d}^{n})]$ is monotonely non-increasing with $n$, then
@@ B.7 Proof of Theorem 3
$\mathbb{E}\left[\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right]=\frac{\sum_{k=1}^{d}\widehat{t}_{k}\tau_{k}}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right).$
@@
$\,\mathbb{E}\left[\widehat{\psi}-\psi\right]$
@@
$\,\left|\mathbb{E}\left[\widehat{\psi}-\psi\right]\right|$
@@
$\operatorname{Var}(\widehat{\psi})=\operatorname{Var}\left(\mathbb{E}\left[\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right]\right)+\mathbb{E}\left[\operatorname{Var}\left(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)\right].$ (19)
@@
$\mathbb{E}\left[\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right]=\frac{\sum_{k=1}^{d}\widehat{t}_{k}\tau_{k}}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)=\frac{\sum_{k=1}^{d}\widehat{t}_{k}(\tau_{k}-\psi)}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)+\psi I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right).$
@@
$\operatorname{Var}\left(\mathbb{E}\left[\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right]\right)\leq\sigma_{n}^{2}+2\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right)+4\sigma_{n}\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right).$
@@
$\mathbb{E}\left[\operatorname{Var}\left(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)\right].$
@@
$\operatorname{Var}(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n})=\frac{I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)}{\left(\sum_{j=1}^{d}\widehat{t}_{j}\right)^{2}}\sum_{k=1}^{d}\widehat{t}_{k}\left(\frac{\mu_{1k}(1-\mu_{1k})}{n\widehat{p}_{k}\widehat{\pi}_{k}}+\frac{\mu_{0k}(1-\mu_{0k})}{n\widehat{p}_{k}(1-\widehat{\pi}_{k})}\right).$
Let $W=\sum_{k=1}^{d}\widehat{t}_{k}\in\sigma(\mathbf{X}^{n},\mathbf{A}^{n})$ be the number of categories that contain both treated and untreated units. On the event $\{W=m\}$ for some $m\in[d]$ we have
$\,\operatorname{Var}(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n})I(W=m)$
$\leq$ $\,\frac{I(W=m)}{4m^{2}}\sum_{k=1}^{d}\widehat{t}_{k}\left(\frac{1}{n\widehat{p}_{k}\widehat{\pi}_{k}}+\frac{1}{n\widehat{p}_{k}(1-\widehat{\pi}_{k})}\right).$
Notice that $\widehat{t}_{k}=I(\widehat{p}_{k}>0,0<\widehat{\pi}_{k}<1)$ and on the event $\{\widehat{p}_{k}>0,0<\widehat{\pi}_{k}<1\}$ we have $n\widehat{p}_{k}\widehat{\pi}_{k}\geq 1,n\widehat{p}_{k}(1-\widehat{\pi}_{k})\geq 1$. We thus have
$\operatorname{Var}(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n})I(W=m)\leq\frac{I(W=m)}{2m^{2}}\sum_{k=1}^{d}\widehat{t}_{k}=\frac{I(W=m)}{2m}.$
Simple observation shows $\operatorname{Var}(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n})I(W=0)=0$. Thus we have
$\,\mathbb{E}\left[\operatorname{Var}\left(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)\right]$
$=$ $\,\sum_{m=1}^{d}\mathbb{E}\left[I(W=m)\operatorname{Var}\left(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)\right]$
$\leq$ $\,\sum_{m=1}^{d}\frac{\mathbb{P}(W=m)}{2m}$
$=$ $\,\frac{1}{2}\mathbb{E}\left[\frac{I(W>0)}{W}\right],$
where the estimated proportion $n\widehat{{\boldsymbol{p}}}\sim\text{Multinomial}(n,{\boldsymbol{p}})$. In the following analysis, we first use Poissonization to reduce the problem to Poisson sampling scheme. Let
$f_{5}(x)=\frac{I(x>0)}{x}+I(x=0)$
and note that $f_{5}$ is non-increasing on $\{0,1,\dots,d\}$. Also as the sample size $n$ increases, the number of levels with both treated and untreated units will not decrease, i.e. $W$ is non-decreasing as we collect more samples in one experiment. This shows
$\mathbb{E}[f_{5}(W)]$
is non-increasing in $n$ and hence by Poissonization Lemma 5 we have
$\mathbb{E}_{n\widehat{{\boldsymbol{p}}}\sim\text{Multi}(n,{\boldsymbol{p}})}[f_{5}(W)]\leq 2\mathbb{E}_{n\widehat{{\boldsymbol{p}}}\sim\text{Poi}(n{\boldsymbol{p}})}[f_{5}(W)]$
where we use $n\widehat{{\boldsymbol{p}}}\sim\text{Poi}(n{\boldsymbol{p}})$ to denote $n\widehat{p}_{i}\sim\text{Poi}(np_{i})$ and $\{n\widehat{p}_{1},\dots,n\widehat{p}_{d}\}$ are independent. This implies
$\,\mathbb{E}\left[\operatorname{Var}\left(\widehat{\psi}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)\right]$
$\leq$ $\,\frac{1}{2}\mathbb{E}_{n\widehat{{\boldsymbol{p}}}\sim\text{Multi}(n,{\boldsymbol{p}})}\left[\frac{I(W>0)}{W}\right]$
$\leq$ $\,\mathbb{E}_{n\widehat{{\boldsymbol{p}}}\sim\text{Poi}(n{\boldsymbol{p}})}\left[\frac{I(W>0)}{W}+I(W=0)\right]$
$\leq$ $\,\mathbb{E}_{n\widehat{{\boldsymbol{p}}}\sim\text{Poi}(n{\boldsymbol{p}})}\left[\frac{I(W>0)}{W}\right]+\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right),$
where the last inequality follows since in the proof of Lemma 1 we show
$\mathbb{P}_{n\widehat{{\boldsymbol{p}}}\sim\text{Poi}(n{\boldsymbol{p}})}(W=0)\leq\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right).$
In the following derivation we will assume $n\widehat{{\boldsymbol{p}}}\sim\text{Poi}(n{\boldsymbol{p}})$ and bound
$\mathbb{E}\left[\frac{I(W>0)}{W}\right].$
Now $W=\sum_{k=1}^{d}I(\widehat{p}_{k}>0,0<\widehat{\pi}_{k}<1)$ is a summation of independent Bernoulli random variable with potentially different mean. Let
$q_{k}=\mathbb{P}(\widehat{p}_{k}>0,0<\widehat{\pi}_{k}<1)=1+\exp(-np_{k})-\exp(-np_{k}(1-\pi_{k}))-\exp(-np_{k}\pi_{k}),$
$\bar{q}=\frac{1}{d}\sum_{k=1}^{d}q_{k}$ and $\overline{W}\sim\text{B}(d,\bar{q})$. We will use the following lemma to relate $W$ with $\overline{W}$.
Lemma 6.
(Theorem 3 in Hoeffding, (1956)) Let $X=\sum_{i=1}^{n}\xi_{i}$ where $\xi_{1},\dots,\xi_{n}$ are independent Bernoulli random variables with parameters $p_{1},\dots,p_{n}$. Let $\bar{p}=\frac{1}{n}\sum_{i=1}^{n}p_{i}$ and $\overline{X}\sim\text{B}(n,\bar{p})$. For any convex function $g$ in the sense that
[…]
we have
$\mathbb{E}[g(X)]\leq\mathbb{E}[g(\overline{X})]$
To use Lemma 6, for any $\delta>0$ let
$f_{\delta}(x)=\frac{I(x>0)}{x}+I(x=0)\left(\frac{3}{2}+\delta\right).$
One can verify $f_{\delta}$ is convex in the sense of Lemma 6 and hence $\mathbb{E}[f_{\delta}(X)]\leq\mathbb{E}[f_{\delta}(\overline{X})]$, i.e.
$\mathbb{E}\left[\frac{I(W>0)}{W}\right]+\left(\frac{3}{2}+\delta\right)\mathbb{P}(W=0)\leq\mathbb{E}\left[\frac{I(\overline{W}>0)}{\overline{W}}\right]+\left(\frac{3}{2}+\delta\right)\mathbb{P}(\overline{W}=0).$
Since $\delta>0$ is arbitrary we conclude
$\mathbb{E}\left[\frac{I(W>0)}{W}\right]+\frac{3}{2}\mathbb{P}(W=0)\leq\mathbb{E}\left[\frac{I(\overline{W}>0)}{\overline{W}}\right]+\frac{3}{2}\mathbb{P}(\overline{W}=0).$
By Lemma 7 we have
$\mathbb{E}\left[\frac{I(\overline{W}>0)}{\overline{W}}\right]\leq\frac{2}{d\bar{q}}$
Since $\pi_{k}\in[\epsilon,1-\epsilon]$, we have
$q_{k}\in[1+\exp(-np_{k})-\exp(-np_{k}(1-\epsilon))-\exp(-np_{k}\epsilon),1+\exp(-np_{k})-2\exp(-np_{k}/2)].$
Note that we assume $p_{k}\geq\alpha/d$ for all $k$ and by using the following two inequalities for $t>0$
$\exp(-t)\leq 1-t+t^{2}/2,$
$\exp(-t)\geq 1-t+t^{2}/2-t^{3}/6$
we have
$q_{k}\geq$  $\,1+\exp\left(-\frac{\alpha n}{d}\right)-\exp\left(-\frac{\epsilon\alpha n}{d}\right)-\exp\left(-\frac{(1-\epsilon)\alpha n}{d}\right)$
$\geq$  $\,1+1-\frac{\alpha n}{d}+\frac{\alpha^{2}n^{2}}{2d^{2}}-\frac{\alpha^{3}n^{3}}{6d^{3}}-1+\frac{\epsilon\alpha n}{d}-\frac{\epsilon^{2}\alpha^{2}n^{2}}{2d^{2}}-1+\frac{(1-\epsilon)\alpha n}{d}-\frac{(1-\epsilon)^{2}\alpha^{2}n^{2}}{2d^{2}}$
$=$ $\,\frac{\alpha^{2}\epsilon(1-\epsilon)n^{2}}{d^{2}}-\frac{\alpha^{3}n^{3}}{6d^{3}}$
$\geq$ $\,\frac{\alpha^{2}\epsilon(1-\epsilon)n^{2}}{2d^{2}}$
in the regime $n\leq\frac{3\epsilon(1-\epsilon)d}{\alpha}$. Thus we have
$\frac{2}{d\bar{q}}\leq\frac{4d}{\alpha^{2}\epsilon(1-\epsilon)n^{2}},$
$\mathbb{P}(\overline{W}=0)=(1-\bar{q})^{d}\leq\left(1-\frac{\alpha^{2}\epsilon(1-\epsilon)n^{2}}{2d^{2}}\right)^{d}\leq\exp\left(-\frac{\alpha^{2}\epsilon(1-\epsilon)n^{2}}{2d}\right).$
Finally we obtain the following bound on $\mathbb{E}\left[\frac{I(W>0)}{W}\right]$ under Poissonization
$\mathbb{E}\left[\frac{I(W>0)}{W}\right]\leq\frac{4d}{\alpha^{2}\epsilon(1-\epsilon)n^{2}}+\frac{3}{2}\exp\left(-\frac{\alpha^{2}\epsilon(1-\epsilon)n^{2}}{2d}\right),$
which yields a bound

### v2

@@ 4 Role of Effect Homogeneity
In this section, we study the role of effect homogeneity in consistent estimation of treatment effects in the regime $n\lesssim d$. Let $\tau_{k}=\mathbb{E}[Y|X=k,A=1]-\mathbb{E}[Y|X=k,A=0]=\mu_{1k}-\mu_{0k}$ be the conditional average treatment effect (CATE) in the $k$-th category. In the high-dimensional regime where $d=d_{n}$ can grow with $n$, $\tau_{k}$’s should be viewed as a triangular array $\{\tau_{nk},1\leq k\leq d_{n},n\geq 1\}$ and we will slightly abuse the notation to denote $\tau_{k}=\tau_{nk}$. Homogeneous effects (i.e. when $\tau_{k}=\psi$ for all $k\in[d]$) can be helpful in terms of estimation. Intuitively, since the treatment effects are the same within each level of covariate, there is no need to consistently estimate CATE $\tau_{k}$ for all possible $d$ categories. One only needs to estimate the CATE within a few categories accurately and due to effect homogeneity, these estimated CATEs generalize to other levels of covariate, which potentially reduces the sample size required for consistent estimation. In this work, we adopt a novel form of approximate effect homogeneity via parameter capturing the extent of heterogeneity. Mathematically, denote
@@
where $\widehat{t}_{k}=\widehat{p}_{k}I(0<\widehat{\pi}_{k}<1)$ is the indicator of whether both treated and untreated samples are observed in the $k$-th level and $\widehat{\tau}_{k}=\widehat{\mu}_{1k}-\widehat{\mu}_{0k}$. The idea is to only estimate the CATE $\tau_{k}$ within those categories with both treated and untreated units and take a weighted average over such categories. We restrict our attention to these categories so that unbiased estimation of $(\mu_{1k},\mu_{0k})$ (and hence $\tau_{k}$) is possible. We again define $/0=0$ 0 if $\sum_{k=1}^{d}\widehat{t}_{k}=0$, i.e. there is no such category that contains both treated and untreated units. Clearly on the event $\left\{\sum_{k=1}^{d}\widehat{t}_{k}=0\right\}$, we cannot obtain any information on treatment effects from $\widehat{\tau}$. Under appropriate scaling conditions, the probability of this “adverse” event converges to 00 even in the case $n\lesssim d$, as summarized in the following lemma.
@@
$\operatorname{Var}(\widehat{\tau})\lesssim\sigma_{n}^{2}+\exp\left(-C^{\prime}(\epsilon)\frac{n^{2}}{n\vee d}\right)+\frac{n\vee d}{n^{2}},$
where $C^{\prime}(\epsilon)$ and the constant behind $\lesssim$” depend on $\epsilon$. As a consequence, the estimator $\widehat{\tau}$ is consistent if $\sigma_{n}\rightarrow 0$ and $d/n^{2}\rightarrow 0$ in the regime $n\lesssim d$.
Theorem 3 shows that if asymptotic effect homogeneity holds (i.e. $\sigma_{n}\rightarrow 0$), then the estimator $\widehat{\tau}$ is consistent in the regime $d/n^{2}\rightarrow 0$. Compared with the rate condition $d/n\rightarrow 0$ required for consistency by the plug-in estimator $\widehat{\psi}$, $\widehat{\tau}$ has a faster convergence rate and enables us to achieve consistency in a wider regime under asymptotic effect homogeneity.
Remark 3.
In the previous version of this manuscript, we consider a slightly different estimator
$\widetilde{\tau}=\frac{\sum_{k=1}^{d}I(\widehat{p}_{k}>0,0<\widehat{\pi}_{k}<1)\widehat{\tau}_{k}}{\sum_{k=1}^{d}I(\widehat{p}_{k}>0,0<\widehat{\pi}_{k}<1)},$
i.e., the estimated CATE $\widehat{\tau}_{k}$ is not weighted using $\widehat{p}_{k}$ and all categories with both treated and untreated units are given the same weight. A similar faster convergence rate is achieved under additional assumption that there exists a constant $\alpha>0$ such that $p_{k}\geq\alpha/d$ for all $k\in[d]$. This assumption is critical to $\widetilde{\tau}$ since it rules out the possibility that $p_{k}$ is very small for some $k\in[d]$, which may induce a large variance in estimating the CATE $\tau_{k}$. On the other hand, the estimator $\widehat{\tau}$ in (9) does not rely on that assumption since we re-weight each category with $\widehat{p}_{k}$ and categories with small proportion will be down-weighted accordingly, hence reducing the overall variance of $\widehat{\tau}$.
@@ 5.1 Average Treatment Effects
The negative result in Theorem 4 can be explained by the fact that the underlying measure of ATE estimation is the product of covariate distribution and propensity score, as shown in analyzing its second-order estimator (Robins et al.,, 2009; Zeng et al.,, 2023). In observational studies where propensity score is unknown, it’s hard to obtain nearly unbiased estimates of $p_{k}\mu_{1k}$ (the summand in $\psi_{1}$) with the observed data. One may write $p_{k}\mu_{1k}=p_{k}\pi_{k}\mu_{1k}/\pi_{k}$ with $p_{k}\pi_{k}\mu_{1k}$ unbiasedly estimable. The extra term $1/\pi_{k}$ could be approximated by a polynomial with possibly diverging degrees
@@ B.6 Proof of Lemma 1
Lemma 5 (Theorem 5.10 in Mitzenmacher and Upfal, (2017)).
Let $\mathbf{X}^{n}\in\mathbb{R}^{d}\sim\text{Multinomial}\left(n,p_{1},\ldots,p_{d}\right)$, $\mathbf{Y}^{n}\in\mathbb{R}^{d}$ and $Y_{i}^{n}\sim\text{Poisson}(np_{i})$ and $Y_{1}^{n},\dots,Y_{d}^{n}$ are independent. Consider a non-negative function $f(x_{1},\dots,x_{d})$, if $\mathbb{E}[f(X_{1}^{n},\dots,X_{d}^{n})]$ is monotonely non-increasing with $n$, then
@@ B.7 Proof of Theorem 3
$\mathbb{E}\left[\widehat{\tau}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right]=\frac{\sum_{k=1}^{d}\widehat{t}_{k}\tau_{k}}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right).$
@@
$\,\mathbb{E}\left[\widehat{\tau}-\psi\right]$
@@
$\,\left|\mathbb{E}\left[\widehat{\tau}-\psi\right]\right|$
@@
$\operatorname{Var}(\widehat{\tau})=\operatorname{Var}\left(\mathbb{E}\left[\widehat{\tau}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right]\right)+\mathbb{E}\left[\operatorname{Var}\left(\widehat{\tau}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)\right].$ (19)
@@
$\mathbb{E}\left[\widehat{\tau}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right]=\frac{\sum_{k=1}^{d}\widehat{t}_{k}\tau_{k}}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)=\frac{\sum_{k=1}^{d}\widehat{t}_{k}(\tau_{k}-\psi)}{\sum_{k=1}^{d}\widehat{t}_{k}}I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)+\psi I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right).$
@@
$\operatorname{Var}\left(\mathbb{E}\left[\widehat{\tau}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right]\right)\leq\sigma_{n}^{2}+2\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right)+4\sigma_{n}\exp\left(-C(\epsilon)\frac{n^{2}}{n\vee d}\right).$ (20)
@@
$\mathbb{E}\left[\operatorname{Var}\left(\widehat{\tau}\mid\mathbf{X}^{n},\mathbf{A}^{n}\right)\right].$
@@
$\operatorname{Var}(\widehat{\tau}\mid\mathbf{X}^{n},\mathbf{A}^{n})=$ $\,\frac{I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)}{\left(\sum_{j=1}^{d}\widehat{t}_{j}\right)^{2}}\sum_{k=1}^{d}\widehat{p}_{k}^{2}I(0<\widehat{\pi}_{k}<1)\left(\frac{\mu_{1k}(1-\mu_{1k})}{n\widehat{p}_{k}\widehat{\pi}_{k}}+\frac{\mu_{0k}(1-\mu_{0k})}{n\widehat{p}_{k}(1-\widehat{\pi}_{k})}\right)$
$\leq$ $\,\frac{1}{4}\frac{I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)}{\left(\sum_{j=1}^{d}\widehat{t}_{j}\right)^{2}}\sum_{k=1}^{d}\widehat{p}_{k}^{2}I(0<\widehat{\pi}_{k}<1)\left(\frac{1}{n\widehat{p}_{k}\widehat{\pi}_{k}}+\frac{1}{n\widehat{p}_{k}(1-\widehat{\pi}_{k})}\right)$
Denote $\mathbf{I}=(I_{1},\dots,I_{d})$ with $I_{k}=I(0<\widehat{\pi}_{k}<1)$ and further note that
$\,\frac{1}{4}\mathbb{E}\left[\frac{I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)}{\left(\sum_{j=1}^{d}\widehat{t}_{j}\right)^{2}}\sum_{k=1}^{d}\frac{\widehat{p}_{k}^{2}I(0<\widehat{\pi}_{k}<1)}{n\widehat{p}_{k}\widehat{\pi}_{k}}\right]$
$=$ $\,\frac{1}{4}\mathbb{E}\left[\frac{I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)}{\left(\sum_{j=1}^{d}\widehat{t}_{j}\right)^{2}}\sum_{k=1}^{d}\widehat{p}_{k}^{2}I(0<\widehat{\pi}_{k}<1)\mathbb{E}\left(\frac{1}{n\widehat{p}_{k}\widehat{\pi}_{k}}\mid\mathbf{X}^{n},\mathbf{I}\right)\right]$
$=$ $\,\frac{1}{4}\mathbb{E}\left[\frac{I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)}{\left(\sum_{j=1}^{d}\widehat{t}_{j}\right)^{2}}\sum_{k=1}^{d}\widehat{p}_{k}^{2}I(0<\widehat{\pi}_{k}<1)\mathbb{E}\left(\frac{1}{n\widehat{p}_{k}\widehat{\pi}_{k}}\mid\mathbf{X}^{n},I_{k}\right)\right],$
where the last equation follows from the independence of treatment assignment within each category after covariates $\mathbf{X}^{n}$ are sampled. It is easy to see from Lemma 6 that for a Binomial random variable $V\sim\text{B}(n,p)$ we have
$\mathbb{E}\left[\frac{1}{V}\mid 0<V<n\right]\leq\frac{2}{(n+1)p\left[1-p^{n}-(1-p)^{n}\right]}.$
Thus we have (assume $n\geq 2$)
$\,\widehat{p}_{k}^{2}I(0<\widehat{\pi}_{k}<1)\mathbb{E}\left(\frac{1}{n\widehat{p}_{k}\widehat{\pi}_{k}}\mid\mathbf{X}^{n},I_{k}\right)$
$\leq$ $\,\frac{2\widehat{p}_{k}^{2}I(0<\widehat{\pi}_{k}<1)}{(n\widehat{p}_{k}+1)\pi_{k}\left[1-\pi_{k}^{n}-(1-\pi_{k})^{n}\right]}$
$\leq$ $\,\frac{2\widehat{p}_{k}^{2}I(0<\widehat{\pi}_{k}<1)}{(n\widehat{p}_{k}+1)\epsilon\left[1-\epsilon^{n}-(1-\epsilon)^{n}\right]}$
$\leq$ $\,\frac{2\widehat{p}_{k}^{2}I(0<\widehat{\pi}_{k}<1)}{(n\widehat{p}_{k}+1)\epsilon\left[1-\epsilon^{2}-(1-\epsilon)^{2}\right]}$
$=$ $\,\frac{\widehat{p}_{k}^{2}I(0<\widehat{\pi}_{k}<1)}{(n\widehat{p}_{k}+1)\epsilon^{2}(1-\epsilon)}$
$\leq$ $\,\frac{\widehat{p}_{k}I(0<\widehat{\pi}_{k}<1)}{\epsilon^{2}(1-\epsilon)n}.$
Sum these terms up, we have
$\frac{1}{4}\mathbb{E}\left[\frac{I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)}{\left(\sum_{j=1}^{d}\widehat{t}_{j}\right)^{2}}\sum_{k=1}^{d}\widehat{p}_{k}^{2}I(0<\widehat{\pi}_{k}<1)\mathbb{E}\left(\frac{1}{n\widehat{p}_{k}\widehat{\pi}_{k}}\mid\mathbf{X}^{n},I_{k}\right)\right]\leq\frac{1}{4\epsilon^{2}(1-\epsilon)}\mathbb{E}\left[\frac{I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)}{n\sum_{j=1}^{d}\widehat{t}_{j}}\right]$
Similarly, we can show
$\frac{1}{4}\mathbb{E}\left[\frac{I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)}{\left(\sum_{j=1}^{d}\widehat{t}_{j}\right)^{2}}\sum_{k=1}^{d}\frac{\widehat{p}_{k}^{2}I(0<\widehat{\pi}_{k}<1)}{n\widehat{p}_{k}(1-\widehat{\pi}_{k})}\right]\leq\frac{1}{4\epsilon(1-\epsilon)^{2}}\mathbb{E}\left[\frac{I\left(\sum_{j=1}^{d}\widehat{t}_{j}>0\right)}{n\sum_{j=1}^{d}\widehat{t}_{j}}\right]$
So we have the following bound on the expected conditional variance: